evaluate limx-0(n/1+n^2+n/4+n^2+n/9+n^2+.....+n/2n^2)


1
Expert's answer
2021-04-19T02:02:08-0400

"lim_{n\rightarrow0}(\dfrac{n}{1+n^2}+\dfrac{n}{4+n^2}+\dfrac{n}{9+n^2}+.....+\dfrac{n}{2n^2})"


The value of all the terms in given limit is 0 except last term


"\Rightarrow lim_{n\rightarrow 0}\dfrac{1}{2n}"


"=\dfrac{1}{0}=\infty"


Hence The given limit does not exist.